OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x2 - 4x + 3.

OpenStudy (anonymous):

when it comes to the parabola questions, it is wise of us to know that the parabola can be described as ax^2+bx+c=0(a is not 0); the x of the vertex is -b/(2a), then we will find out the value y: doing like this: a=1;b=-4 so the vertex P(x1,y1) x1=-b/(2a)=4/2=2 y1=2^2-4*2+3=-1 so P(2,-1)